OpenStudy (anonymous):

Determine if triangle DEF with coordinates D (3, 2), E (4, 6), and F (7, 3) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral?

OpenStudy (whpalmer4):

Do you know the distance formula for the distance between two points \((x_1,y_1),(x_2,y_2)\)? \[d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\] what do you know about the lengths of the sides of an equilateral triangle?

OpenStudy (anonymous):

They all must be equal..

OpenStudy (anonymous):

I used GeoGebra and saw that it isn't a equilateral triangle

OpenStudy (anonymous):

Just having trouble of how to make it a equilateral triangle

OpenStudy (whpalmer4):

Okay, are two of the sides equal in length?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Still there? @whpalmer4

OpenStudy (whpalmer4):

well, then take a point which isn't equidistant from two of the other points, and move it so that it is :-)

OpenStudy (anonymous):

can you give me an example?

OpenStudy (whpalmer4):

okay, what is the length of the side that is equal for two sides?

OpenStudy (anonymous):

4.12

OpenStudy (whpalmer4):

no, it's not. it might be a number close to that. stick to exact numbers, which means if you have a square root that you can't simplify, write \(\sqrt{17}\) (or whatever) instead.

OpenStudy (whpalmer4):

in fact, you're usually best off avoiding decimals entirely until you have the final answer, and even then, use a decimal only if the problem asks for one.

OpenStudy (whpalmer4):

Okay, so your problem is to find a point which is \(\sqrt{17}\) away from two of the other points (which are \(\sqrt{17}\) apart from each other).

OpenStudy (whpalmer4):

You can do this with the distance formula (which is really the Pythagorean theorem in disguise). What are the two points which are \(\sqrt{17}\) apart and that you will keep for your new triangle?

OpenStudy (anonymous):

i honestly don't know..

OpenStudy (whpalmer4):

you told me you found the distance between the points. did you, or didn't you?

OpenStudy (whpalmer4):

you have 3 points. 2 pairs of those points are sqrt(17) apart. 1 pair is not. pick one of the pairs that is, and tell me what the points are.

OpenStudy (anonymous):

(3,2)

OpenStudy (whpalmer4):

okay, now give me one of the other pairs, which is sqrt(17) away from (3,2)

OpenStudy (anonymous):

(4,6)

OpenStudy (whpalmer4):

Here's a question for you: do you think the problem is asking for you to find the exact points to use to make this an equilateral triangle, or just describe in general terms how to make an equilateral triangle?

OpenStudy (anonymous):

I already have the exacts points etc. i just need to find out points that do make a equilateral triangle..

OpenStudy (whpalmer4):

you have the exact points to make this an equilateral triangle? how do you know that without knowing what your equilateral triangle is?

OpenStudy (anonymous):

I'm trying to figure out how to create a equilateral triangle

OpenStudy (anonymous):

since coordinates D (3, 2), E (4, 6), and F (7, 3) is not q equilateral triangle i must find points that do make one

OpenStudy (anonymous):

@sourwing